OpenStudy (anonymous):

Using the equation y equals x plus eleven, what is the value of y when x is twelve? y = x + 11 x y 2 13 5 16 12 ? 30 41 seventeen eleven twenty-one twenty-three

OpenStudy (anonymous):

y = x+11 so when x= 12 y= 12+11 = 23 twenty three

OpenStudy (anonymous):

thank you